How many mL of water must be added to 250mL of 0.578 M CaCl2 solution to give a 0.349 M solution?



1
Expert's answer
2022-02-24T16:42:01-0500

CaCl2+H2O =CaO +2HCL

Moles of CaCl2

=250×0.578=144.5 moles

Mole ratio is 1:1 so moles of water =144.5


Ml=144.5/0.349=414.04ML